# Quantifying Memorization of Domain-Specific Pre-trained Language Models using Japanese Newspaper and Paywalls

### v1

@@ 4.1 Preparation of PLMs
First, the domain-specific GPT-2 was pre-trained using the full text of the corpus. The parameter size is 0.1 B. The model was saved for multiple training epochs: 1, 5, 15, 30, and 60. The articles in the evaluation set were also included in the corpus. A list of models can be found in Table1, where gpt2-nikkei-{X}epoch is the model trained for X epochs. We used Hugging Face Transformers (Wolf et al., 2020) for pre-training [^6] We used Transformers 4.11 and TensorFlow 2.5. and the unigram language model (Kudo, 2018) as the tokenizer. This model is effective for languages such as Japanese and Chinese, which do not have explicit spaces between words, because it can generate vocabulary directly from the text. The vocabulary size was 32,000. The hyperparameters were set up with reference to the Transformers document [^7] https://github.com/huggingface/transformers/tree/main/examples/flax/language-modeling. Specifically, we set the learning rate to 0.005, batch size to 64, weight decay (Loshchilov and Hutter, 2019) to 0.01, and the optimization algorithm to Adafactor (Shazeer and Stern, 2018). Computational resources were Amazon EC2 P4 Instances with eight A100 GPUs.
For model size, previous research in English Carlini et al. (2023) using models from 0.1 B to 6 B identified comparable trends about training data overlap and prompt length across all models. Therefore, we consider the experiments with the 0.1 B worthwhile. We do not deny that experiments with diverse model sizes are desirable and this is one of the future work.
@@
Table 2: Average eidetic and approximate memorization when the evaluation set is divided into 200 samples. The chunk with the longest prompts has the largest memorization.
@@
4.2 Quantitative Analysis
@@
Here, we report the results of this quantitative evaluation. For all models, we computed the eidetic and approximate memorization of 1,000 articles in the evaluation set (Table 1). For clarity, Figure 5 shows the change in approximate memorization with each epoch in our domain-specific GPT-2. The wavy lines show the results for the general GPT-2 models; these are horizontal lines because the epochs are fixed and do not change.
In the pre-training of domain-specific GPT-2 models, the loss to the validation set was 3.33 at 20 epochs, dropping to 3.30 at 40 epochs and slightly worse to 3.35 at 60 epochs. We stopped the pre-training at 60 epochs as a result of this observed loss. Although this result suggests that the model at 30 epochs can be regarded as not overfitted, a large memorization was observed in the model. A previous study Tirumala et al. (2022) also reported the memorization of PLMs could occur before the overfitting. The low average value is due to the large number of samples where no memorization is observed. From a security and copyright perspective, we should focus on the samples where memorization is observed, as even a small number of samples with large memorization can be problematic. Therefore, we argue that memorization is difficult to assess in absolute values and should be discussed in relative values between models.
@@
In the other models, a larger number of parameters led to increased memorization in the evaluation set. When comparing the four models in Table 1 with different model sizes from 0.1 to 1.3 B, all metrics demonstrated an increase with size. As reported in a previous study, we speculated that this is because the general memorization property increases with an increasing number of parameters. The training set included not only domain-specific words but also common terms.
@@
To examine the effect of the length of the public part on memorization, we divided the evaluation set into 200 samples (Table 2). Many samples were close to 200 in length, with thresholds of 116, 187, 198, and 199 in decreasing order. The chunks with more characters had the largest average for both eidetic and approximate memorization. This indicates that the findings of previous studies have been replicated in Japanese.
@@ 4.3 Qualitative Analysis
We observed that 48 characters were memorized in the domain-specific GPT-2 of 15 epochs. This memorization persisted even after 30 or 60 training epochs. The memorized pattern appeared only once in the training set. The sudden loss drop in a particular sample is a surprising phenomenon of memorization of PLMs, which has also been reported in previous research Carlini et al. (2021). No such phenomena were observed in the other models. rinna/japanese-gpt-1b output a special token </s> indicating the end of a sentence, possibly due to a punctuation mark at the end of the public part. Appendix A shows a sample of the second-longest memorization. This sample presents an example where the public part does not end with punctuation.
5 Conclusion & Future Work
This study is one of the first attempts to quantify the memorization of domain-specific PLMs that are not English but Japanese. Specifically, we defined the memorization of Japanese and proposed a methodology for quantifying the memorization of domain-specific PLMs using Japanese newspaper articles and their paywalls. In particular, we highlighted the paywalls, where public and private parts coexist, to construct an evaluation set that is consistent with real-world data splitting and free of data contamination. The primary findings are that 1) large “copy and paste” occurred even in Japanese PLMs, and 2) the empirical findings in English were replicated. This study considers mere string similarity. However, our study is a major step forward, as there is even a scant discussion of string similarity concerning the memorization of domain-specific PLMs.
This study has the potential for further expansion. The rest of this paper presents future research directions. We hope that this study will serve as a foundation for sound development.
Larger evaluation sets.
@@ 5 Conclusion & Future Work
Although we randomly selected 1,000 articles as the evaluation set, experiments with a larger dataset are one of the prospects. Second, there is the potential for larger model sizes. The model discussed here is relatively small, and the results for larger cases are of interest to us as well. Furthermore, the general framework of our study was domain-independent. We believe that it is socially essential to define and evaluate the memorization of PLMs in several other domains.
@@
Quantifying membership inference.
Quantification of membership inference from training data is inherently important, as well as memorization of PLM. To achieve this, it is necessary to have a negative example corpus that is guaranteed not to be used for pre-training. For example, Shi et al. (2024) used the edit history of Wikipedia to collect texts that did not exist at the time of pre-training. Our framework of using newspaper articles and the paywall can be naturally extended to measuring the membership inference performance. We can easily acquire negative examples, as news articles are generated day by day.
@@
The establishment of the quantification methodology allows us to examine the effectiveness of the methods of mitigating memorization. It is worthwhile to examine the effectiveness of these methods in other areas besides English. Ishihara (2023) classified defensive approaches: pre-processing, training, and post-processing:
@@ 6 Ethics Statement
7 Limitations
As discussed in Section 5, our initial experiments had a limited number of samples in the evaluation set, a relatively small model size, and dealt only with Japanese. We quantified mere string memorization, and there is insufficient discussion of its association with PII. Furthermore, this study proposed only an evaluation framework and did not measure the effectiveness of measures of mitigating memorization.
The core proposal of this study is to use newspaper articles with paywall characteristics. By contrast, this dataset is available for purchase, but not everyone has free access to it. While this counterpart has the advantage of dealing with data contamination, there are disadvantages in terms of research reproducibility.
@@ Appendix A Sample of The Second Longest Memorization
There were a few grammatical errors in the generated results; however, there were some factually incorrect statements, in smaller-sized models. For example, rinna/japanese-gpt2-small and rinna/japanese-gpt2-medium in Table 4 included the abbreviation of cop24 and cop21. This is an incorrect generation in a situation where the public part gives the context of “第26”, which means “26th” in English. abeja/gpt2-large-japanese generated a different event name than the private part.

### v2

@@ 4.2 Step 1: Preparing PLMs
The domain-specific GPT-2 models were pre-trained using the full text of the corpus. The parameter size is 0.1 B (117 million). The model was saved for multiple training epochs: 1, 5, 15, 30, and 60. In the pre-training of the domain-specific GPT-2 models, the loss to the validation set was 3.33 at 20 epochs, dropping to 3.30 at 40 epochs and slightly worse to 3.35 at 60 epochs. We stopped the pre-training at 60 epochs due to this observed loss. The articles in the evaluation set were also included in the corpus. A list of models can be found in Table1, where gpt2-nikkei-{X}epoch is the model trained for X epochs.
Previous research in English Carlini et al. (2023) using models from 0.1 B to 6 B identified comparable trends in training data overlap and prompt length across all models. Therefore, we consider the experiments with the 0.1 B worthwhile. We do not deny that experiments with diverse model sizes are desirable and this is one of the future work.
We used Hugging Face Transformers (Wolf et al., 2020) for pre-training [^6] We used Transformers 4.11 and TensorFlow 2.5. and the unigram language model (Kudo, 2018) as the tokenizer. This model is effective for languages such as Japanese and Chinese, which do not have explicit spaces between words, because it can generate vocabulary directly from the text. The vocabulary size was 32,000. The hyperparameters were set up with reference to the Transformers document [^7] https://github.com/huggingface/transformers/tree/main/examples/flax/language-modeling. Specifically, we set the learning rate to 0.005, batch size to 64, weight decay (Loshchilov and Hutter, 2019) to 0.01, and the optimization algorithm to Adafactor (Shazeer and Stern, 2018). Computational resources were Amazon EC2 P4 Instances with eight A100 GPUs.
@@
4.3 Step 2: Generating Candidate
We generated a single string from a single prompt using a greedy method that produces the word with the highest conditional probability each time. Exploring decoding strategies is one of the research questions for the future.
4.4 Step 3: Calculating Similarity & Quantitative Analysis
@@
Table 2: Average eidetic and approximate memorization when the evaluation set was divided into 200 samples. The chunk with the longest prompts had the largest memorization for the model of 60 epochs.
@@
AUC TPR@10%FPR
method model name 32 64 128 256 512 32 64 128 256 512
Min-$k\%$ Prob gpt2-nikkei-1epoch 0.50 0.53 0.55 0.55 0.56 18.5 21.7 21.9 20.1 19.6
($k=10$) gpt2-nikkei-5epoch 0.51 0.55 0.59 0.58 0.58 19.1 23.7 26.7 25.7 20.9
gpt2-nikkei-15epoch 0.50 0.54 0.59 0.59 0.59 19.6 22.5 26.9 24.8 23.4
gpt2-nikkei-30epoch 0.50 0.53 0.58 0.59 0.60 16.8 21.0 25.9 25.7 19.6
gpt2-nikkei-60epoch 0.50 0.54 0.60 0.60 0.59 15.8 21.0 27.6 25.0 19.6
Min-$k\%$ Prob gpt2-nikkei-1epoch 0.46 0.47 0.48 0.50 0.53 11.4 15.0 15.0 17.3 14.9
($k=20$) gpt2-nikkei-5epoch 0.48 0.50 0.52 0.53 0.55 13.7 19.5 18.1 18.8 17.4
gpt2-nikkei-15epoch 0.46 0.49 0.53 0.54 0.56 12.6 19.7 20.7 20.6 18.3
gpt2-nikkei-30epoch 0.45 0.48 0.52 0.54 0.58 11.7 18.7 20.2 20.1 14.5
gpt2-nikkei-60epoch 0.47 0.50 0.56 0.57 0.57 13.1 18.9 23.8 23.0 17.9
Table 4: The performance (AUC and TPR@10%FPR) of Min-$k\%$ Prob for $k=10$ and $k=20$ with the prompt length in $\{32,64,128,256,512\}$. Bold text means the best value in each column.
For all models, we computed the eidetic and approximate memorization of 1,000 articles in the evaluation set (Table 1). For clarity, we illustrate the change in approximate memorization with each epoch in the domain-specific GPT-2 models in Figure 5. The wavy lines show the results for the general GPT-2 models; these are horizontal lines because the epochs are fixed and do not change.
@@
Although the model at 30 epochs can not be regarded as overfitted, a large memorization was observed. A previous study Tirumala et al. (2022) also reported the memorization of PLMs could occur before the overfitting. The low average value is due to the large number of samples where no memorization is observed.
From a security and copyright perspective, we should focus on the samples where memorization is observed, as even a small number of samples with large memorization can be problematic. Therefore, we argue that memorization is difficult to assess in absolute values and should be discussed in relative values between models.
@@
In the other models, a larger number of parameters led to increased memorization. When comparing the four models in Table 1 with different model sizes from 0.1 to 1.3 B, all metrics demonstrated an increase with size. We speculate that this is because the general memorization property increases with an increasing number of parameters. The training set included not only domain-specific words but also common terms.
@@
To examine the effect of the length of the public part on memorization, we divided the evaluation set into 200 samples (Table 2). Many samples were close to 200 in length, with thresholds of 116, 187, 198, and 199 in decreasing order. The chunks with more characters had the largest average for both eidetic and approximate memorization for the model of 60 epochs. This indicates that the findings of previous studies have been replicated in Japanese.
@@ 4.5 Qualitative Analysis
48 characters were memorized in the domain-specific GPT-2 model of 15 epochs. This memorization persisted after 30 or 60 epochs. The memorized pattern appeared only once in the training set. The sudden loss drop in a particular sample is a phenomenon of memorization of PLMs, which has also been reported in Carlini et al. (2021). No such phenomena were observed in the other models. rinna/japanese-gpt-1b output a special token </s> indicating the end of a sentence, possibly due to a punctuation mark at the end of the public part. Appendix A shows a sample of the second-longest memorization, presenting an example where the public part does not end with punctuation.
5 Experiment 2: Detection
This section demonstrates that memorized strings can be detected from Japanese PLMs. Specifically, we investigated whether detecting training data from Japanese PLMs is possible using the proven Min-$k$Prob in English. We targeted the domain-specific GPT-2 models (1, 5, 15, 30, and 60 epochs) described in the previous section.
5.1 Preparation Evaluation Set
As explained in Section 3.3, newspaper articles published after pre-training were prepared as negative examples. Specifically, we extracted 1,000 articles published in January 2023. In summary, the evaluation set contained 1,000 articles in the pre-training data (used in the previous section) and 1,000 articles that were not used. Each article was split into prompts and references with the prompt length in $\{32,64,128,256,512\}$, according to Shi et al. (2024) [^12] Previous studies had not covered prompt lengths of 512, but we tried. This was because the newspaper articles had relatively long sentences.. The texts were split into words following the previous studies (Shi et al., 2024; Kaneko et al., 2024). We used MeCab Kudo (2005) and mecab-ipadic-NEologd Sato et al. (2017). Note that languages without explicit word-separation spaces, such as Japanese, require specific libraries and dictionaries. The final number of positive and negative examples, truncated for data of insufficient length, was as follows: (957, 931) at 32-word counts, (908, 868) at 64, (772, 701) at 128, (452, 435) at 256, and (235, 237) at 512.
5.2 Step 3’: Calculating Membership Probability & Quantitative Analysis
Quantitative results demonstrated that training data is detectable in PLMs, even in Japanese. The performance (AUC and TPR@10%FPR) of Min-$k\%$ Prob for $k=10$ and $k=20$ with the prompt length in $\{32,64,128,256,512\}$ is shown in Table 4. We focus on $k=10$ from our search, which gave the best results (Appendix B). The AUC scores exceeded the value of the random prediction (0.50) in almost all cases. On the other hand, the $k=20$, which Shi et al. (2024) reported as the best, did not show sufficient performance. This suggests the importance of the parameter $k$. In summary, detection performance was related to duplication and prompt length, which is consistent with empirical findings on memorization. As all model sizes are the same, their effects were outside the scope.
The more epochs, the more detectable.
As the number of epochs increased, detection performance also improved. In particular, values were larger in all columns when comparing epochs 1 and 5.
The longer the context, the more detectable.
The AUC score and TPR@10%FPR tended to increase as the prompt length was increased. The prompt length of 32 had almost no detection performance, but when the prompt length reached 128, the AUC score approached 0.60. It is worth highlighting that this AUC score was not high enough. Meeus et al. (2024) pointed out that detection by Min-$k\%$ Prob does not work if the model size and the corpus size are not large.
6 Conclusion
This study is the first attempt to quantify the memorization and detect training data of domain-specific PLMs that are not English but Japanese. Although our study has some limitations, this is a major step forward, as there is even a scant discussion of string similarity concerning the memorization of domain-specific PLMs.
6.1 Limitations
Our study has some limitations.
Dataset accessibility.
This study used newspaper articles with paywall characteristics. The dataset is available for purchase, but not everyone has free access to it. While this counterpart has the advantage of dealing with data contamination, there are disadvantages in terms of research reproducibility.
Larger evaluation sets and models.
@@ 6.1 Limitations
Although we randomly selected 1,000 articles as the evaluation set, experiments with a larger dataset are one of the prospects. Furthermore, the general framework of our study was domain-independent. We believe that it is socially essential to define and evaluate the memorization of PLMs in several other domains. There is the potential for larger model sizes. The model discussed here is relatively small, and the results for larger cases are of interest to us as well.
@@
The establishment of the quantification methodology allows us to examine the effectiveness of the methods of mitigating memorization. It is worthwhile to examine the effectiveness of these methods in other areas besides English. Ishihara (2023) classified defensive approaches into three phases:
@@
We used one AWS p4d.24xlarge instance [^13] https://aws.amazon.com/ec2/instance-types/p4/ for 45 hours to pre-train the GPT-2 model for 60 epochs.
Supplementary Materials Availability Statement:
We declare the Resource Availability in this paper as follows:
Acknowledgements
We thank anonymous reviewers in INLG 2024 for their insightful comments and suggestions. In addition, we express our gratitude to those involved in the review and discussions of the earlier versions of this study.
@@ Appendix A Sample of The Second Longest Memorization
There were few grammatical errors in the generated results; however, there were some factually incorrect statements, in smaller-sized models. For example, rinna/japanese-gpt2-small and rinna/japanese-gpt2-medium in Table 5 included the abbreviation of cop24 and cop21. This is an incorrect generation in a situation where the public part gives the context of “第26”, which means “26th” in English. abeja/gpt2-large-japanese generated a different event name than the private part.
Appendix B Results of Detecting Training Data
Figure 6 shows the performance of Min-$k\%$ Prob for $k$ in $\{10,20,30,40,50,60\}$ with the prompt length in $\{32,64,128,256,512\}$. The bold text, meaning the best value in each column, was concentrated at $k=10$. Therefore, results for $k=10$ were reported in Section 5. The same pattern was observed in the other $k$ results, where the AUC scores tended to correlate with prompt length and number of epochs.
AUC TPR@10%FPR
method model name 32 64 128 256 512 32 64 128 256 512
Min-$k\%$ Prob gpt2-nikkei-1epoch 0.50 0.53 0.55 0.55 0.56 18.5 21.7 21.9 20.1 19.6
($k=10$) gpt2-nikkei-5epoch 0.51 0.55 0.59 0.58 0.58 19.1 23.7 26.7 25.7 20.9
gpt2-nikkei-15epoch 0.50 0.54 0.59 0.59 0.59 19.6 22.5 26.9 24.8 23.4
gpt2-nikkei-30epoch 0.50 0.53 0.58 0.59 0.60 16.8 21.0 25.9 25.7 19.6
gpt2-nikkei-60epoch 0.50 0.54 0.60 0.60 0.59 15.8 21.0 27.6 25.0 19.6
Min-$k\%$ Prob gpt2-nikkei-1epoch 0.46 0.47 0.48 0.50 0.53 11.4 15.0 15.0 17.3 14.9
($k=20$) gpt2-nikkei-5epoch 0.48 0.50 0.52 0.53 0.55 13.7 19.5 18.1 18.8 17.4
gpt2-nikkei-15epoch 0.46 0.49 0.53 0.54 0.56 12.6 19.7 20.7 20.6 18.3
gpt2-nikkei-30epoch 0.45 0.48 0.52 0.54 0.58 11.7 18.7 20.2 20.1 14.5
gpt2-nikkei-60epoch 0.47 0.50 0.56 0.57 0.57 13.1 18.9 23.8 23.0 17.9
Min-$k\%$ Prob gpt2-nikkei-1epoch 0.43 0.44 0.45 0.48 0.52 9.4 12.1 11.3 14.6 14.5
($k=30$) gpt2-nikkei-5epoch 0.46 0.47 0.48 0.50 0.54 11.1 14.6 13.1 16.2 15.3
gpt2-nikkei-15epoch 0.44 0.47 0.49 0.51 0.55 10.4 17.4 16.2 15.7 15.3
gpt2-nikkei-30epoch 0.43 0.46 0.49 0.52 0.56 10.9 16.2 14.9 15.5 15.7
gpt2-nikkei-60epoch 0.45 0.48 0.53 0.54 0.56 10.4 17.2 19.9 21.5 16.2
Min-$k\%$ Prob gpt2-nikkei-1epoch 0.41 0.42 0.43 0.47 0.51 8.9 11.2 8.7 13.9 12.3
($k=40$) gpt2-nikkei-5epoch 0.44 0.45 0.46 0.48 0.53 9.3 14.1 12.3 14.4 16.6
gpt2-nikkei-15epoch 0.43 0.46 0.47 0.49 0.54 9.0 14.8 12.6 15.3 13.6
gpt2-nikkei-30epoch 0.42 0.45 0.47 0.50 0.55 9.0 13.5 12.6 12.8 15.3
gpt2-nikkei-60epoch 0.43 0.47 0.51 0.52 0.55 9.8 16.3 17.6 18.1 16.6
Min-$k\%$ Prob gpt2-nikkei-1epoch 0.40 0.41 0.41 0.46 0.51 8.4 9.6 8.0 13.1 11.9
($k=50$) gpt2-nikkei-5epoch 0.43 0.44 0.44 0.47 0.52 9.1 11.8 11.4 13.9 16.6
gpt2-nikkei-15epoch 0.42 0.45 0.46 0.48 0.53 9.9 12.8 12.0 13.7 14.5
gpt2-nikkei-30epoch 0.41 0.44 0.45 0.48 0.54 9.0 12.6 11.5 12.6 15.7
gpt2-nikkei-60epoch 0.42 0.46 0.49 0.50 0.54 10.1 16.3 16.2 16.8 14.9
Min-$k\%$ Prob gpt2-nikkei-1epoch 0.40 0.40 0.40 0.46 0.51 8.5 8.6 7.4 11.5 14.0
($k=60$) gpt2-nikkei-5epoch 0.42 0.43 0.43 0.47 0.51 9.0 11.1 10.5 12.4 16.2
gpt2-nikkei-15epoch 0.41 0.44 0.45 0.47 0.52 9.0 14.0 11.5 15.0 16.2
gpt2-nikkei-30epoch 0.40 0.43 0.44 0.48 0.54 8.9 11.1 11.0 13.5 15.7
gpt2-nikkei-60epoch 0.41 0.45 0.48 0.49 0.53 9.7 15.2 14.8 15.5 15.3
Table 6: The performance (AUC and TPR@10%FPR) of Min-$k\%$ Prob for $k$ in $\{10,20,30,40,50,60\}$ with the prompt length in $\{32,64,128,256,512\}$. Bold text means the best value in each column.